This is on a take home assingment.
Question: Which equation describes an allipse?
(A) -3x^2-7x-3y^2+2y-9=0
(B) 2x^2-7x+4y^2-9y=-13
(C) 5y^2-7y-12x^2-3x+2=0
(D) 2y^2-7x-9y=-1
Thank you very much!!!!!!!!!!!!!! Answer by mukhopadhyay(490) 
 Answer is C.
5y^2-7y-12x^2-3x+2=0
=>5(y-7/10)^2-12(x+3/24)^2=-2+49/20+9/288 (Right side is +ve, it is an ellipse)
-3x^2-7x-3y^2+2y-9=0
=>3x^2+7x+3y^2-2y=-9
=>3(X+7/6)^2+3(Y-2/6)^2=-9+49/12+4/12 (Right side is -ve, so cannot be ellipse)
..
2x^2-7x+4y^2-9y=-13
=>2(x-7/4)^2+4(y-9/8)^2=-13+49/8+81/16 (is -ve, so cannot be ellipse)
..
The last equation represents parabola